# Improving accuracy of tree-tensor network approach by optimization of network structure

### v1

@@ III.3 Richardson Model
The result that the stochastic schemes do not work well for the Richardson model also manifests itself in the TTN structure. Figure 15 shows the TTN structure obtained by a calculation of type (2D) with $\chi_{\rm opt}=40$ for $J_{\rm R}=1$ and $N=128$, which yielded a relatively high energy. In this TTN, the spins around $i=N/2$ are scattered into various branches, and the TTN should be unsuitable for expressing the entanglement among those spins. The reason why the stochastic scheme obtained such a TTN may be understood as follows. Roughly speaking, in stochastic schemes, the TTN structure is randomized by probabilistic selection of local TTN structures at the early stage of the calculation where the effective temperature is high, which scatters the spins around $i=N/2$ into various branches. Then, the TTN structure is optimized as the effective temperature is lowered. However, in the randomized TTN, the spins near $i=N/2$ are surrounded by nearly classical-ordered spins at the sites away from $i=N/2$. In the process of local reconnection of isometries in the structural optimization, the position swap of the blocks consisting of such classically ordered spins does not change the EE or truncated rDM weight at the center bond very much. As a result, the “slope” of the cost function of EE or truncated rDM weight in the TTN-structural space becomes small in such a TTN, which slows down the structural optimization and makes it easier to be trapped even in a shallow local minimum. This situation may not change even if $J_{\rm R}$ becomes larger, as long as the profile of $\langle S^{z}_{i}\rangle$ is flat and the effect of local reconnection of isometries on the cost function is small. Thus, the poor accuracy of the stochastic schemes may be attributed to the fact that those schemes create an effective initial TTN structure where the slope of cost function in the local structural update is small and the optimization flow slows down and even freezes.
@@
To confirm the above scenario, we performed four runs of the calculations of non-stochastic schemes [schemes (B) and (C)] with $\chi_{\rm opt}=40$, starting from a PBT in which the bare spins are arranged randomly. The relative error in the variational energy, $\delta e$, obtained are shown in Figs. 10 and 11 (by downward triangles). The results indicate that the accuracy of the calculation is as poor as that of stochastic schemes [schemes (D) and (E)], supporting the above scenario that the stochastic optimization is roughly equivalent to the non-stochstic one starting from a randomized TTN. These observations underscore the importance of appropriately selecting the initial TTN as well as the optimization scheme according to the system treated.
@@
In non-stochastic schemes (B) and (C), there are some cases where the relative error in the variational energy is large, as seen in Figs. 10 and 11. This may also be caused by the phenomenon that once the structural optimization flow strays into the TTN structure with a small slope of the cost function, it is stuck there.
@@ IV Concluding Remarks
In summary, we have evaluated the performance of the TTN structural optimization algorithm for variational calculations of quantum many-body systems. We have examined several types of the algorithm, which are specified by the initial TTN structure, the scheme to update the local connection of isometries in the structural optimization, and the bond dimention at which the structural optimization is performed. We applied them to the random XY-exchange model under random magnetic field and the Richardson model and analyzed the variational energies and the EEs on the auxiliary bonds. The results for the random XY model, on the one hand, demonstrate that the TTN structural optimization algorithm works effectively to improve the accuracy of the calculation for the model. In particular, we have found that stochastic schemes, which probabilistically select the local connection of isometries in the optimization process, yield the best results in the random average of the variational energy, regardless of the initial TTN structure. For the Richardson model, on the other hand, the results show that the structural optimization algorithm of an appropriately selected type can realize a lower variational energy than those obtained by the calculation without TTN structural optimization. Nevertheless, we have also found that, particularly in the calculations of stochastic schemes, the TTN structure obtained by the optimization can stuck in a structure that is not suitable for representing the lowest energy state of interest, resulting in poor calculation accuracy. This indicates the importance of appropriately choosing the initial TTN structure as well as the schemes of selecting local connections of isometries in the structural optimization, depending on the model.
@@
The TTN structural optimization algorithm we proposed searches for the optimal structure through local reconnections of the network. Therefore, it is essential to devise a way to realize fast and steady convergence to the optimal structure. We have tested the stochastic schemes according to Eqs. (5) and (6). For the random XY model, the stochastic schemes succeeded in obtaining the lowest variational energies independent of the initial TTN structure, while non-stochastic schemes yielded results that are good but not the best and depend on the initial TTN. In the random XY model, the landscape of variational energy in the TTN-structure space is expected to be complex, and the optimization flow of the non-stochastic scheme is likely to be trapped in local minima, leading to an increase in the variational energy. The stochastic scheme should be effective in promoting the escape from such local minima. In contrast, for the Richardson model, the stochastic schemes yielded high variational energies, while the non-stochastic schemes achieved relatively good results depending on the initial TTN. A possible explanation for this result is that once the TTN structure is randomized at the early stage of stochastic optimization, the slope of EE in the TTN structural space almost vanishes, and the optimization flow can be easily trapped in a shallow minimum. In order to avoid the problem, it may be effective to try multiple calculations from various initial TTNs as well as to adjust the annealing schedule in the stochastic optimization. The idea of the replica exchange method[36, 37], which has recently been applied to the structural optimization of tensor networks with loops[38], is also expected to be promising.
@@
While we focused in this work on the performance of the TTN structural optimization algorithm in the variational calculation of quantum many-body systems, it has many other potential applications. One promising application is the one to an efficient representation of a given high-rank tensor in the TTN format. In the algorithm we proposed, each step involves the diagonalization of the effective Hamiltonian expanded in the truncated Hilbert space of the central area to obtain the ground-state wave function (process 3 of Table 1). By replacing this process with the one to obtain a wave function that has a maximum fidelity to the given high-rank tensor, one can optimize only the TTN structure without changing the TTN state itself.[39] Such an application is expected to be useful for machine learning[40, 41] and quantics tensor-train approach[42, 43, 44], to which tensor networks have been applied actively in recent years. An application of a similar algorithm using bond mutual information as a cost function to the Born machine has been explored very recently.[45]
@@ Appendix A Preparation of initial TTN
In our calculation, isometries in the initial TTN are prepared as follows. For example, let us consider the MPN shown in Fig. 1 (a). First, we focus on the subsystem consisting of the leftmost two spins at the sites $i=1$ and $2$ in the MPN and construct its block Hamiltonian. Then, we generate the leftmost isometry using the $\chi$ eigenvectors of the block Hamiltonian with the $\chi$-lowest eigenenergies. If the number of eigenvectors of the block Hamiltonian does not exceed $\chi$, all eigenvectors are adopted in the isometry. Next, we focus on the subsystem consisting of the leftmost two-site block and the neighboring spin at the site $i=3$, construct and diagonalize the block Hamiltonian of the subsystem, and generate the new isometry using the $\chi$-lowest-energy eigenvectors. We iterate these procedures from the left edge to the center of the MPN, where the canonical center is located, and perform the same operation from the right edge to the center to eventually prepare the initial MPN. (To be precise, we generate only the isometries in the initial TTN and not the singular-value matrix, the latter not necessary for the start of our calculation.) The other initial TTNs are also prepared in the same way, with the only difference being that we change the order of merging the spin blocks. In the case of PBT, the spin blocks are merged from the boundary of PBT, where the bare spins are located, to the “top” of PBT, where the canonical center is, so that PBT shown in Fig. 1 (b) is constructed. In the case of the tSDRG, in each step of the preparation procedure, we construct the interaction Hamiltonians of all the pairs of spin blocks, diagonalize them, and calculate the maximum level spacing $\Delta_{\rm max}^{\rm I}$ of their energy spectrum. Then, among all the spin-block pairs, we select the one with the largest $\Delta_{\rm max}^{\rm I}$ to be merged at the step. See Ref. [29] for the details.
@@
In general, in our algorithm, the structure of the initial TTN can be arbitrary. The closer the structure of the initial TTN is to the optimal one, the faster the convergence of the calculation is. As for the initial isometries, any isometries can be used as long as they satisfy the orthonormal condition Eq. (1). Preparing the isometries using random vectors satisfying the condition may be a reasonable choice. It is also possible to use a product state, a TTN with $\chi=1$, as the initial TTN. However, in that case, it is necessary to take measures against the problem of the appearance of an isometry whose three legs have the dimension unity, mentioned in Sec. II, since the TTN with $\chi=1$ of a generic structure inevitably contains such an isometry and the calculation will be stuck. A possible measure to avoid the problem is employing a MPN as the structure of the initial TTN so that the central area has at least two physical bonds, whose dimension is usually larger than unity, as the incoming bonds, and increasing the dimension of auxiliary bonds sufficiently before starting the structural optimization.
@@
The problem of the isometry having three legs of the dimension unity may also occur accidentally during the TTN optimization calculation. In our calculations of the present work, we encountered the problem in several cases of the calculation of the types where the initial TTN was (2) PBT for the Richardson model with $(J_{\rm R},N)=(64,64)$, if we did not take the exceptional treatment mentioned below. The problem occurred when the initial PBT included isometries whose outgoing bond was composed of the $\chi$-lowest-energy eigenstates of the block Hamiltonian all of which belonged to the subspace of the same magnetization. Since the states of the spin block represented by such an isometry can take only the single magnetization in the truncated Hilbert space, the XY-exchange interactions between the block and the rest of the system become zero, and the block is decoupled. Then, if two or more such blocks are connected to the central area at a step in the TTN calculation, the rank of the ground-state wave function in one (or more) singular-value decomposition becomes unity, generating a bond with dimension $\chi=1$. The cumulation of such a process may result in the emergence of the isometry with all three legs having the dimension $\chi=1$.
@@ Appendix B Details of calculations
In our TTN calculations, we computed physical quantities such as variational energies as a function of the upper bound of the bond dimension, $\chi$, with increasing $\chi$. The calculations for the random XY model and the Richardson model were performed in the following manner.
@@
random XY model: For the calculation without structural optimization [scheme (A)], we performed the variational calculation to optimize the isometries with increasing $\chi$ as $\chi=16(100),24(20),32(20),40(20)$, where the figures in the parentheses are the maximum number of sweeps iterated at each $\chi$. If the variational energy converged within the range of $10^{-5}$ before the maximum number of sweeps was reached, we finished the calculation at the value of $\chi$ and proceeded to the next $\chi$. We then adopted the lowest variational energy obtained during the sweeps as the variational energy for each $\chi$.
@@
For the calculations with the structural optimization [schemes (B)-(E)], we first performed the calculations with the structural-optimization procedure with $\chi=\chi_{\rm opt}$; two cases of calculations with $\chi_{\rm opt}=16$ and $32$ were carried out. We iterated the structural-optimization calculations until the TTN structure and the variational energy converged, and then adopted the TTN structure obtained as the optimized one in the following calculation. If the calculation did not converge within $80$ sweeps, the TTN structure obtained after the last sweep was used as the optimized one. After the optimized TTN structure was determined, we performed the variational calculation to optimize the isometries with increasing $\chi$ as $\chi=16,24,32,40$ for $\chi_{\rm opt}=16$ and $\chi=32,40$ for $\chi_{\rm opt}=32$. The calculation at each $\chi$ was continued until the variational energy converged or the number of sweeps reached $20$.
@@
Richardson model: The calculations for the Richardson model were performed basically in the same manner as for the random XY model, with slight differences (see below).
@@
For the calculation without structural optimization, we performed the calculation to optimize the isometries with increasing $\chi$ as $\chi=20(100),40(20),60(20),80(20)$, where the figures in the parentheses denote the number of sweeps iterated at each $\chi$. Here, differently from the case of the random XY model, we did not monitor the convergence of the variational energy but instead performed the calculation of the stated number of sweeps. We then adopted the lowest variational energy obtained during the sweeps as the variational energy at the value of $\chi$. For the calculations with the structural optimization, we performed the calculations with the structural-optimization procedure with $\chi=\chi_{\rm opt}=20$ or $40$ until the convergence of the TTN structure and the variational energy (within the range of $10^{-6}$) was achieved or the number of sweeps reached $80$. After that, we performed the variational calculation to optimize the isometries with increasing $\chi$ as $\chi=20,40,60,80$ for $\chi_{\rm opt}=20$ and $\chi=40,60,80$ for $\chi_{\rm opt}=40$. We iterated the calculations of 20 sweeps for each $\chi$ to obtain the variational energy as a function of $\chi$.
Appendix C Exact solution of Richardson model

### v2

@@ III.3 Richardson Model
The result that the stochastic optimization does not work well for the Richardson model also manifests itself in the TTN structure. Figure 15 shows the TTN structure obtained by the stochastic optimization for the initial PBT with $\chi_{\rm opt}=40$ for $J_{\rm R}=1$ and $N=128$. The TTN of this structure yielded the relative error $\delta e=5\times 10^{-9}$ ($2\times 10^{-8}$) for $\chi=80$ ($60$), which seems quite small but $7\times 10^{2}$ ($2\times 10^{3}$) times larger than the error by the optimal TTN. In the TTN shown in Fig. 15, the spins around $i=N/2$ are scattered into various branches; This TTN should fail in expressing the entanglement among those spins. The reason why the stochastic optimization generated such a TTN may be understood as follows. Roughly speaking, in the stochastic optimization, the TTN structure is randomized by probabilistic selection of local TTN structures at the early stage of the calculation where the effective temperature is high, which scatters the spins around $i=N/2$ into various branches. Then, the TTN structure is optimized as the effective temperature is lowered. However, in the randomized TTN, the spins near $i=N/2$ are surrounded by nearly classical-ordered spins at the sites away from $i=N/2$. In the process of local reconnection of isometries, the position swap of the clusters of such classically ordered spins does not change the EE or the truncated rDM weight at the center bond very much. As a result, the “slope” of the cost function of the bond EE or the truncated rDM weight in the TTN-structural space becomes small. In such a TTN, the structural optimization slows down and can be trapped even in a shallow local minimum. This situation may not change even if $J_{\rm R}$ becomes larger, as long as the profile of $\langle S^{z}_{i}\rangle$ away from $i=N/2$ is flat and the effect of local reconnection of isometries on the cost function is small.
@@
To confirm the above scenario, we have performed four runs of the non-stochastic structural optimization with $\chi_{\rm opt}=40$, starting from PBTs with random shuffling of bare spin sites. The relative errors in the variational energy, $\delta e$, thus obtained are shown in Figs. 10 and 11 (by downward triangles). The results indicate that the accuracy of the calculation is as poor as that of the stochastic optimization, supporting the above scenario that the stochastic optimization is roughly equivalent to the non-stochastic one starting from a randomized TTN. These observations underscore the importance of appropriately selecting the initial TTN as well as the optimization scheme according to the system treated.
@@
As for the non-stochastic optimizations, there are some cases where $\delta e^{(^{\prime})}$ is relatively large, as seen in Figs. 10 and 11. We anticipate that this may also be understood by the scenario that once the structural optimization strays into the TTN structure with a small slope of the cost function, it is stuck there.
@@ IV Concluding Remarks
In summary, we have evaluated the performance of the TTN structural optimization algorithm for variational calculations of quantum many-body systems. We have examined several types of optimization, which are specified by the initial TTN structure, the scheme to update the local connection of isometries, and the bond dimension at which the structural optimization is performed. We applied them to the random XY-exchange model under random magnetic fields and the Richardson model and analyzed the variational energies and the EEs on the auxiliary bonds. On the one hand, the results for the random XY model demonstrate that the TTN structural optimization algorithm works effectively to improve the accuracy of the calculation for the model. In particular, the stochastic optimization, which probabilistically selects the local connection of isometries in the optimization process, yields the best results in the random average of the variational energy, regardless of the initial TTN structure. On the other hand, the results for the Richardson model show that the structural optimization algorithm of an appropriately selected type can realize a lower variational energy than those obtained by the calculation without structural optimization. Nevertheless, we have also found that, particularly in the calculations of the stochastic optimization, the TTN structure obtained by the optimization can get stuck in a structure that is not suitable for representing the lowest-energy state of interest, resulting in poor accuracy.
@@
The TTN structural optimization algorithm we proposed searches for the optimal structure through local reconnections of the network. Therefore, it is essential to devise a way to realize fast and steady convergence to the optimal structure. We have tested the stochastic optimization scheme according to Eqs. (5). [See also Appendix C for the test of another stochastic optimization using the probability of Eq. (6).] For the random XY model, the stochastic optimization succeeded in obtaining the lowest variational energies independent of the initial TTN structure, while the non-stochastic optimizations yielded results that are good but not the best and depend on the initial TTN. In the random XY model, the landscape of the variational energy in the TTN-structure space is expected to be complex, and the optimization of the non-stochastic scheme is likely to be trapped in a local minimum, leading to poor accuracy. The stochastic optimization should be effective in promoting the escape from such a local minimum. In contrast, for the Richardson model, the stochastic optimization has yielded high variational energies, while the non-stochastic optimizations have achieved relatively good results depending on the initial TTN. A possible explanation for this result is that once the TTN structure is randomized at the early stage of the stochastic optimization, the slope of the bond EE in the TTN structural space almost vanishes, and the optimization can be easily trapped in a shallow minimum.
The results of the present study indicate the importance of appropriately choosing the initial TTN structure and the scheme of selecting local connections of isometries in the structural optimization, depending on the model. Generally speaking, for systems with randomness, it is naturally expected that the stochastic optimization is essential. In contrast, for systems without randomness, the non-stochastic optimization should be the first choice since the stochastic scheme involves the risk that the structural optimization gets stuck in a randomized TTN structure. Then, in either case, it must be effective to try the optimizations of various types as many as possible. As observed in Sec III, the TTN structure that yields a good variational energy at a small $\chi$ usually produces a good energy also at a large $\chi$ (see Figs. 4 and 9). Therefore, we can test several types of optimization calculations at a small $\chi$ where the computational cost is negligible, and then adopt the TTN structure with the best precision, which is expected to achieve the best precision also for a large $\chi$. Another promising solution may be to employ a scheme based on the idea of the replica exchange method[37, 38], which has recently been applied to the structural optimization of tensor networks with loops[39].
We have compared the accuracy of various types of structural optimizations at the same upper bound of the bond dimensions, $\chi$, focusing on the ability of TTNs constrained by the same $\chi$ to represent quantum states. Here, it is also important to consider the computational costs to conduct the calculations. In general, the computational cost of the calculation using a generic TTN is larger than that using MPN since isometries in MPN always have at least one physical bond with dimension $d$, which is often smaller than $\chi$, while isometries in TTN have no such constraint. The memory required for storing isometries scales as $\mathcal{O}(N\chi^{3})$ for generic (binary) TTN and $\mathcal{O}(N\chi^{2}d)$ for MPN. The computational costs to renormalize the spin and Hamiltonian operators are respectively $\mathcal{O}(N^{2}\chi^{4})$ and $\mathcal{O}(N\chi^{5})$ for TTN whereas $\mathcal{O}(N^{2}\chi^{3}d)$ and $\mathcal{O}(N\chi^{3}d^{2})$ for MPN. The cost to perform the update of isometries in the variational calculation depends on the algorithm; For the two-tensor update algorithm, the cost scales as $\mathcal{O}(NN_{\rm int}\chi^{5})$ for TTN and $\mathcal{O}(NN_{\rm int}\chi^{3}d^{2})$ for MPN, where $N_{\rm int}$ is a coefficient depending on the number of terms in the Hamiltonian. For the single-tensor update, the cost is $\mathcal{O}(NN_{\rm int}\chi^{4})$ for TTN and $\mathcal{O}(NN_{\rm int}\chi^{3}d)$ for MPN. Considering these differences, it may be a reasonable option to take advantage of the small cost of MPN, depending on the purpose and situation. Alternatively, if one wishes to utilize the high capability of the optimal TTN for representing a quantum state, it should be effective to first perform the structural optimization of TTN at a small $\chi$ as mentioned above, and then improve isometries for a large $\chi$ by the single-tensor update algorithm, whose computational cost is relatively small.
The structural optimization based on the least-EE principle has been proven to be powerful not only for TTNs but also for MPNs[22] and the tensor networks of entanglement renormalization.[39] Meanwhile, it has been shown recently that for PEPS, the structural optimization based on the principle of selecting the network geometry that maximizes the bond EE is effective for improving the accuracy.[40] Although this maximum-EE principle for PEPS seems incompatible with the least-EE principle, these two are consistent with each other as follows. In TTNs, each bond corresponds to a bipartition of the system that occurs when the bond is cut. The least-EE principle minimizes the appearance of bonds corresponding to a bipartition with a large EE. However, when a physical spin is strongly entangled with other spins, the TTN must represent the large entanglement precisely. Consequently, in a TTN optimized according to the least-EE principle, the strongly entangled spins are connected by as few isometries as possible in the lower layers (regions near the boundary) of the network. In such a TTN, the bonds with large EE that are inevitable for representing the target state accurately are confined to small regions in lower layers, thereby minimizing the bond EEs in the whole TTN. In fact, our calculation of the structural optimization for the Richardson model has yielded the TTN structure in which the spins entangled strongly are located close to each other. Similar TTNs have also been obtained for the hierarchical chain[16] and the Rainbow chain.[30] On the other hand, the optimization principle for PEPS that maximizes the bond EEs yields the network structure in which the strongly entangled spins are connected by a small number of bonds (directly by a single bond, if possible), making the distance between those strongly entangled spins in the PEPS network as small as possible. It is thereby understood that the maximum-EE principle for PEPS realizes the situation occurring in the lower layer of the TTN obtained by the least-EE principle. We note that a similar strategy to connect the strongly entangled spins at the lower layer of the TTN has also been employed in the entanglement-based tSDRG.[15]
@@
While we have focused on the performance of the TTN structural optimization algorithm in the variational calculation of quantum many-body systems, it has many other potential applications. A promising application is an efficient representation of a given high-rank tensor in the TTN format. In the algorithm we proposed, each step involves the diagonalization of the effective Hamiltonian expanded in the truncated Hilbert space of the central area to obtain the ground-state wave function (process 3 of Table 1). By replacing this process with the one to obtain a wave function that has a maximum fidelity to the given high-rank tensor, one can optimize only the TTN structure without changing the TTN state itself.[41] Such an application is expected to be useful for machine learning[42, 43] and quantics tensor-train approach[44, 45, 46], to which tensor networks have been applied actively in recent years. An application of a similar algorithm using bond mutual information as a cost function to the Born machine has been explored very recently.[47]
@@ Acknowledgements.
The data that support the findings of this article are openly available in [48].
@@ Appendix A Preparation of initial TTN
In our calculation, isometries in the initial TTN are prepared as follows. For example, let us consider the MPN shown in Fig. 1 (a). First, we focus on the subsystem consisting of the leftmost two spins at the sites $i=1$ and $2$ in the MPN and construct its block Hamiltonian. Then, we generate the leftmost isometry using the $\chi$ eigenvectors of the block Hamiltonian with the $\chi$-lowest eigenenergies. If the number of eigenvectors of the block Hamiltonian does not exceed $\chi$, all eigenvectors are adopted in the isometry. Next, we focus on the subsystem consisting of the leftmost two-site block and the neighboring spin at the site $i=3$, construct and diagonalize the block Hamiltonian of the subsystem, and generate the new isometry using the $\chi$-lowest-energy eigenvectors. We iterate these procedures from the left edge to the center of the MPN, where the canonical center is located, and perform the same operation from the right edge to the center to eventually prepare the initial MPN. (To be precise, we prepare only the isometries in the initial TTN and do not generate the singular-value matrix, which is not necessary for the start of our calculation.) The other initial TTNs are also prepared in the same way, with the only difference that we change the order of merging the spin blocks. In the case of PBT, the spin blocks are merged from the boundary of PBT, where the bare spins are located, to the “top” of PBT, where the canonical center is, so that PBT shown in Fig. 1 (b) is constructed. In the case of the tSDRG, in each step of the preparation procedure, we construct the interaction Hamiltonians of all the pairs of spin blocks, diagonalize them, and calculate the maximum level spacing $\Delta_{\rm max}^{\rm I}$ of their energy spectrum. Then, among all the spin-block pairs, we select the one with the largest $\Delta_{\rm max}^{\rm I}$ to be merged at the step. See Ref. [29] for the details.
@@
In general, in our algorithm, the structure of the initial TTN can be arbitrary. The closer the structure of the initial TTN is to the optimal one, the faster the convergence of the calculation is. As for the initial isometries, any isometries can be used as long as they satisfy the orthonormal condition Eq. (1). Preparing the isometries using random vectors satisfying the condition may be a reasonable choice. It is also possible to use a product state, a TTN with $\chi=1$, as the initial TTN. However, in that case, it is necessary to take measures against the problem of the appearance of an isometry whose three legs have the dimension unity, mentioned in Sec. II, since the TTN with $\chi=1$ of a generic structure inevitably contains such an isometry and the calculation will be stuck. A possible measure to avoid the problem is employing an MPN as the initial TTN so that the central area has at least two physical bonds, whose dimension is usually larger than unity, and increasing the dimension of auxiliary bonds sufficiently before starting the structural optimization.
@@
The problem of the isometry having three legs of dimension unity may also occur accidentally during the TTN optimization calculation. In the present work, we encountered the problem in several cases of the calculations where the initial TTN was PBT for the Richardson model with $(J_{\rm R},N)=(64,64)$, if we did not take the exceptional treatment mentioned below. The problem occurred when the initial PBT included isometries whose outgoing bond was composed only of the bases having the same magnetization. Since the states of the spin block represented by such an isometry can take only the single magnetization in the truncated Hilbert space, the XY-exchange interactions between the block and the rest of the system become zero, and the block is decoupled. Then, if two or more such blocks are connected to the central area, the rank of the ground-state wave function in one (or more) singular-value decomposition becomes unity, generating a bond with dimension $\chi=1$. The cumulation of such a process may result in the emergence of the isometry with all three legs having the dimension $\chi=1$.
@@ Appendix B Details of calculations
In our TTN calculations, we computed physical quantities such as variational energies as a function of the upper bound of the bond dimension, $\chi$, by increasing $\chi$. The calculations for the random XY model and the Richardson model were performed in the following manner.
@@
random XY model: For the calculation without structural optimization, we performed the variational calculation to optimize the isometries by increasing $\chi$ as $\chi=16(100),24(20),32(20),40(20)$, where the figures in the parentheses are the maximum number of sweeps iterated at each $\chi$. If the maximum number of sweeps was reached or the variational energy converged within the range of $10^{-5}$, we finished the calculation at the value of $\chi$ and proceeded to the next $\chi$. We then adopted the lowest variational energy obtained during the sweeps at each $\chi$ as the variational energy for the value of $\chi$.
@@
For the calculations with the structural optimization, we first performed the calculations with the structural optimization procedure at $\chi=\chi_{\rm opt}$; Two cases of calculations with $\chi_{\rm opt}=16$ and $32$ were carried out. We iterated the structural optimization calculations until the TTN structure and the variational energy converged, and then adopted the TTN structure obtained as the optimized one in the following calculation. If the calculation did not converge within $80$ sweeps, the TTN structure obtained after the last sweep was used as the optimized one. After the optimized TTN structure was determined, we performed the variational calculation to optimize the isometries by increasing $\chi$ as $\chi=16,24,32,40$ for $\chi_{\rm opt}=16$ and $\chi=32,40$ for $\chi_{\rm opt}=32$. The calculation at each $\chi$ was continued until the variational energy converged or the number of sweeps reached $20$.
@@
Richardson model: The calculations for the Richardson model were performed in a manner similar to those for the random XY model, with slight differences (see below).
@@
For the calculation without structural optimization, we performed the calculation to optimize the isometries by increasing $\chi$ as $\chi=20(100),40(20),60(20),80(20)$, where the figures in the parentheses denote the number of sweeps iterated at each $\chi$. Here, differently from the case of the random XY model, we did not monitor the convergence of the variational energy but instead performed the calculation of the stated number of sweeps. We then adopted the lowest variational energy obtained during the sweeps as the variational energy at the value of $\chi$. For the calculations with the structural optimization, we performed the calculations with the structural optimization procedure at $\chi=\chi_{\rm opt}=20$ or $40$ until the convergence of the TTN structure and the variational energy (within the range of $10^{-6}$) was achieved or the number of sweeps reached $80$. After that, we performed the variational calculation to optimize the isometries by increasing $\chi$ as $\chi=20,40,60,80$ for $\chi_{\rm opt}=20$ and $\chi=40,60,80$ for $\chi_{\rm opt}=40$. We iterated the calculations of 20 sweeps for each $\chi$ to obtain the variational energy as a function of $\chi$.
Appendix C Results of stochastic calculation using the probability $P^{(2)}$ in Eq. (6)
Figure 16: Random averages of the relative reduction in the variational energy, $\delta r(\chi)]$, for the random XY model obtained by the calculations where the initial TTN is (a) MPN, (b) PBT, and (c) tSDRG-TTN. $P^{(1)}$” and $P^{(2)}$” represent respectively the stochastic optimizations according to the probability $P^{(1)}$ [Eq. (5)] and $P^{(2)}$ [Eq. (6)]. The results for the optimization using the truncated rDM weight (“rDMwgt”) are also presented for comparison.
Figure 17: Relative error in the variational energy, $\delta e(\chi)$, for the Richardson model obtained by the calculations where the initial TTN is (a) MPN and (b) PBT. The model parameters are $J_{\rm R}=1$ and $N=128$. $P^{(1)}$” and $P^{(2)}$” represent respectively the stochastic optimizations according to the probability $P^{(1)}$ [Eq. (5)] and $P^{(2)}$ [Eq. (6)]. Among the four runs, the result of the run yielding the lowest variational energy at $\chi=80$ is plotted. The results for the optimization using the truncated rDM weight (“rDMwgt”) are also presented for comparison.
In the stochastic optimization of the TTN structure discussed in the main text, we have performed the probabilistic selection of the local connection of the isometries according to the probability $P^{(1)}$ in Eq. (5). As mentioned in Secs. II and III.1, it is also possible to adopt the probability $P^{(2)}$ in Eq. (6) for the probabilistic selection, instead of $P^{(1)}$. We have also performed the calculations of the stochastic optimization with $P^{(2)}$ for all cases where the calculations with $P^{(1)}$ were carried out, for both the random XY model and the Richardson model. We have then compared the performance of those calculations.
Figures 16 and 17 show the results of the relative reduction $\delta r(\chi)]$ and the relative error $\delta e(\chi)$ of the variational energy for the random XY model and the Richardson model, respectively, obtained by the stochastic optimizations with $P^{(1)}$ and $P^{(2)}$. (Note that the data of the calculations using $P^{(1)}$ are the same as those shown in Figs. 4 and 9 in the main text.) We find that both optimization schemes achieve almost the same accuracy for those models: while the results of the optimization using $P^{(1)}$ may be slightly more accurate than those using $P^{(2)}$, the differences are minor compared to the differences from the results of the other structural optimization schemes. In addition, we have analyzed the average values, maximum values, and the histograms of EEs on the auxiliary bonds, and confirmed that the stochastic optimizations with $P^{(1)}$ and $P^{(2)}$ yield essentially the same results. Therefore, we conclude that the stochastic optimizations using $P^{(1)}$ and $P^{(2)}$ exhibit nearly equivalent performance, at least for the models treated in the present study. In practice, it may be sufficient to perform the calculation with $P^{(1)}$, and there is no particular need to execute the additional calculation with $P^{(2)}$.
Appendix D Exact solution of Richardson model